I found a problem in a text book, but I wanted to know if I was correct in determining that there was no solution to the problem.
The problem is the cuber root of X=x-6; the book I found stated that the answer was 8, but I have tried every approach to solve this problem and maybe I am missing a step.
I tried to take the difference of two cubes and set the equation equal to zero, but that didn't work. I also expanded the equation and set it equal to zero, but that didn't work either. I don't know how they came up with the answer 8, but I hope you can help.

Answer by stanbon(75887)

Have you tried graphing the two sides of the equation?
Graph y=x^(1/3) and y=x-6 on the same set of axis and
find the point of intersection. It will probably be
at x=8.
Cheers,
Stan H.